25 g of CH3OH dissolved in 45 g of C6H5OH . calculate molality of solution if solution contains high amount of volume is 1L then also calculat normality and molarity of solution

Grade:12

1 Answers

Jit
6 years ago
Mole of ch3oh=25/32molality= 25*1000/(32*45)Volm.of soln. =1 mlMolarity=25/(32*1)Normality=molarity*n-factor=25/(32*1)*1